OpenStudy (anonymous):

What is 0.0314 written in scientific notation? A. 3.14 × 10–3 B. 3.14 × 103 C. 3.14 × 102 D. 3.14 × 10–2

OpenStudy (solomonzelman):

move decimal right 2 decimals places

OpenStudy (ranga):

In scientific notation, the first number has to be between 1 and 10. So 0.0314 = 3.14 /100 = 3.14 / 10^2 = ?